OpenStudy (anonymous):

A bag contains 4 white, 3 blue, and 5 red marbles Find the probability of choosing a blue marble, then a red marble if the marbles are not replaced

OpenStudy (anonymous):

@Loser66 @LegendarySadist

OpenStudy (anonymous):

12

OpenStudy (anonymous):

3

OpenStudy (anonymous):

A 5/44 b 15/35 c 2/3 d 1/5

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

A 5/44 b 15/35 c 2/3 d 1/5

OpenStudy (anonymous):

wow just wanted help if you cant help then don't be rude

OpenStudy (pink.1d):

wait i almost solved it

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

A 5/44 b 15/35 c 2/3 d 1/5

OpenStudy (anonymous):

these are the chiocis to chose from

OpenStudy (pink.1d):

5/44

OpenStudy (pink.1d):

A

OpenStudy (anonymous):

are u positive or u just guessing

OpenStudy (pink.1d):

positive i just solved it y wud i guess/??

OpenStudy (pink.1d):

r u taking a test?

OpenStudy (anonymous):

yes will you help me with the last question plz

OpenStudy (pink.1d):

i can explain if u want

OpenStudy (pink.1d):

ill check it out :) could u check my recent question too maybe you can help

OpenStudy (anonymous):

using the same info A bag contains 4 white, 3 blue, and 5 red marbles Find the probability of choosing 2 white marbles in a row if the marbles are not replaced.

OpenStudy (pink.1d):

can u medal me? im working on it

OpenStudy (pink.1d):

ok so 4/12*3/11 = 0.0909090909 convert that to fraction

OpenStudy (pink.1d):

or just 12/132 = 6/66 = 1/11

OpenStudy (anonymous):

a 1/11 b 1/9 c 2/3 d 1/16

OpenStudy (pink.1d):

its a

OpenStudy (anonymous):

yup wbu

OpenStudy (pink.1d):

can u fan me back pleasE? i fanned u

OpenStudy (anonymous):

oh and thanks for the help

OpenStudy (anonymous):

yeah it is the end of the school year so I want to get a good grade

OpenStudy (anonymous):

ok thanks agine bye

OpenStudy (pink.1d):

np :)